# Lin Li1,2∗,  Wei Chen1∗,  Jiahui Li3,  Kwang-Ting Cheng1,2  Long Chen1†

### v2

@@ 4.1 Experimental Setup
Table 2: Performance(%) comparison with various GSR models on the SWiG [10] dataset.
@@
ISL [10] ${}_{\text{ECCV'20}}$ Close-ended 39.36 30.09 18.62 22.73 7.72
JSL [10] ${}_{\text{ECCV'20}}$ 39.94 31.44 18.87 24.86 9.66
GSRTR [11] ${}_{\text{BMVC'21}}$ 40.63 32.15 19.28 25.49 10.10
CoFormer [36] ${}_{\text{CVPR'22}}$ 44.66 35.98 22.22 29.05 12.21
SituFormer [35] ${}_{\text{AAAI'22}}$ 44.20 35.24 21.86 29.22 13.41
GSRFormer [37] ACM ${}_{\text{ACM MM'22}}$ 46.53 37 .48 23.32 31.53 14.23
OpenSU [47] ${}_{\text{ICCVW'23}}$ Open-ended 50.10 41.20 26.56 34.27 15.70
@@
4.2 Comparison with State-of-the-Art Methods
We compared our Relation-R1 for unified relation comprehension with the existing state-of-the-art binary relation detection and N-ary relation detection approaches in Table 1 and Table 2, respectively.
[…]
4.3 Diagnostic Experiment
Table 3: Performance comparison (%) across various cognitive CoT strategies trained on single binary relation detection and $N$-ary relation detection, separately. Blue indicates metrics without correct verb constraints.
  Binary Relation $N$-ary Relation
@@ 4.3 Diagnostic Experiment
Figure 3: The statistics of binary relation rewards.
Figure 4: The statistics of $N$-ary relation rewards.
Reward Analysis. We visualized the reward variation curves for binary and N-ary relation detection in Figure 3 and Figure 4, respectively. It can be seen that at the beginning stage, both tasks can obtain decent rewards due to the foundation laid by SFT. With the gradual progress of GRPO, the rewards of both tasks show a gradually increasing trend. In binary relation detection, guided with progressive CoT yields a slightly enhanced overall reward trajectory, although the improvement is not significant. This might be since the gap between the CoT generated by MLLM (e.g., Qwen [39]) and that produced by the template is not obvious. Conversely, N-ary relation detection exhibits a significant surge in reward, primarily attributed to the model’s emergent capability to generate synonymous grounded situation frames (c.f. §5).
@@
Figure 5: The statistics of binary relation completion lengths.
Figure 6: The statistics of $N$-ary relation completion lengths.
@@
The evolution of completion length during GRPO training is depicted in Figure 5 and Figure 6. Consistent with the reward trends, the completion lengths for binary relation detection exhibit a gradual increase as GRPO progresses, suggesting the generation of more detailed and descriptive outputs over time. Notably, the template-based CoT strategy yields significantly shorter completions compared to the more flexible MLLM-generated and progressive CoT strategies, since only the cognitive process is involved. The progressive CoT guidance, by enabling the model to explore synonymous relational expressions and incorporate more nuanced details, leads to an increase in the average completion length.
Table 4: Few-Shot GSR performance comparison.
@@
GSRTR [11] ${}_{\text{BMVC'21}}$ One-Shot 0.35 0.10 0.01 0.08 0.00
CoFormer [36] ${}_{\text{CVPR'22}}$ 0.63 0.23 0.02 0.14 0.01
@@
$w$/$o$ RL 27.31 10.21 1.43 9.35 1.22
GSRTR [11] ${}_{\text{BMVC'21}}$ 4-Shot 6.95 4.32 1.83 2.79 0.65
CoFormer [36] ${}_{\text{CVPR'22}}$ 12.39 8.27 3.93 5.28 1.25
@@
$w$/$o$ RL 34.88 17.14 3.94 15.63 3.38
Few-Shot N-ary Relation Detection. We also evaluated Relation-R1’s few-shot generalization capability (Table 4). In the one-shot setting, Relation-R1 achieved dramatic improvements over prior work, reaching 31.67% (Verb) and 3.41% (Grnd-all) compared to CoFormer’s [36] 0.63% and 0.01%, respectively. This advantage is maintained in the 4-shot setting, where Relation-R1 reached 38.02% (Verb) and 5.92% (Grnd-all), significantly surpassing previous methods. These results clearly demonstrate Relation-R1’s superior ability to generalize and perform robust N-ary relation detection, setting a new SOTA for few-shot GSR.
Joint Binary and N-ary Relation Detection.
Table 5: Performance comparison (%) across various cognitive CoT strategies trained on joint binary relation detection and $N$-ary relation detection, separately. Blue indicates metrics without correct verb constraints.
  Binary Relation $N$-ary Relation
@@
Beyond training Relation-R1 solely on binary or N-ary relation detection tasks (as discussed in §3), we also investigated the joint training of both tasks within the same Relation-R1 model. The results of this joint training are presented in Table 5333The Grnd-all metric excludes ground-truth bounding boxes represented as [-1, -1, -1, -1] for calculation.. From the analysis of these results, three primary conclusions can be drawn: 1) When utilizing both the SFT and RL strategies, our Relation-R1 consistently achieves higher performance across nearly all metrics for both binary and N-ary relation detection tasks. Similar to single-task training, the MLLM-generated CoT strategy underperforms in the N-ary task. This is likely due to the inherent generality of the CoT generated by Qwen [39], which may not precisely align with the ground-truth annotations, consequently leading to lower Verb accuracy. 2) Our proposed progressive CoT-guided paradigm consistently achieves the highest performance across all evaluated metrics for both tasks in the joint learning setting. 3) The joint progressive CoT-guided training across both binary and N-ary tasks fosters a synergistic mutual promotion effect. This is evidenced by the notable gains observed: for instance, a Recall of 27.19% in Table 5 vs. 22.57% in Table 3 (for binary relation detection), and a Verb accuracy of 73.25% in Table 5 vs. 71.04% in Table 3 (for N-ary relation detection).
5 Qualitative Results
@@
Figure 7: Performance(%) comparison between ASMv2 and Relation-R1 over all predicates on test set of PSG [17] dataset. The orange area denotes the predicate distribution of the training set.
Figure 8: Qualitative results of binary relation detection.
Figure 9: Qualitative results of N-ary relation detection.
Figure 10: Qualitative results of synonymous in N-ary relation detection.
@@
Comparison over All Predicates. We visualized predicate-level comparison against ASMv2 [8] and the predicate distribution in Figure 7. As seen, Relation-R1 outperforms ASMv2 by a clear margin on both prevalent and rare predicates. These results clearly validated the effectiveness and efficiency of our approach for the SGG task in the open-ended setting.
@@
Binary and N-ary Relation Detection Results. We visualized some examples of binary and N-ary relation detection in Figure 8 and Figure 9, respectively. As seen, for binary relation detection (c.f. Figure 8), the visualized thinking process reveals how Relation-R1 effectively decomposes the complex task into constituent cognitive steps: object classification, localization, and relation inference. For instance, Relation-R1 can first focus on the main objects (e.g., “cat” and “train”) and then infer their positions and relationships. Similarly, for N-ary relation detection (c.f. Figure 9), Relation-R1 showcases its understanding by first identifying the core action (crowning) and then methodically grounding each associated semantic role – Agent (woman), Object (tiara), and Recipient (woman) – to its corresponding entity and bounding box. This structured analysis underscores Relation-R1’s ability to parse complex events involving multiple participants and their distinct roles within the scene.
@@
Synonymous Visualization. Figure 10 displays synonymous results observed in N-ary relation detection. As evinced by these results, our progressive CoT-guided training paradigm empowers the model to generate diverse synonymous expressions, encompassing entities, semantic roles, and activities. For instance, in Figure 10(top), Relation-R1 demonstrates the ability to output synonymous agents such as “person” and “man”. Furthermore, the model can generate semantically similar and plausible activities, exemplified by “barbecues” and “grills” (c.f. Figure 10(bottom)). Even within the same activity “rests”, it can differentiate and utilize varying semantic role names like “item” and “bodypart” (c.f. Figure 10(row 2)). These synonymous outputs indicate a more flexible and robust semantic understanding beyond rigid lexical matching, thereby significantly enhancing the model’s generalization capabilities to diverse linguistic formulations and real-world variations.
6 Conclusion
This paper reveals critical limitations in MLLM visual relation understanding, particularly for N-ary relationships and language-induced bias. We propose Relation-R1, a unified framework with a two-stage approach: establishing foundational reasoning and structured outputs via cognitive CoT-guided SFT, followed by RL with GRPO for refinement, promoting visual-semantic grounding and robust generalization. Furthermore, progressively applied cognitive CoT guidance, transitioning from template-based to MLLM-generated reasoning, further enhances generalization, especially for synonymous N-ary relations. Extensive experiments on PSG and SWiG validate Relation-R1’s effectiveness, demonstrating SOTA performance in both binary and N-ary relation detection. Integrating progressive CoT-guided SFT with R1-inspired RL refinement, Relation-R1 advances MLLMs towards more human-like relational reasoning.
Appendix A Implementation Details
We utilized Qwen2.5-VL-3B [39] as our base model. Relation-R1 was trained on an 8xA100 80GB GPU server. During the SFT phase, we employed a total batch size of 32, with 4 samples processed per training step. For the GRPO component, we used a total batch size of 48. The learning rates were set to 2e-5 for SFT and 1e-6 for GRPO, respectively. Qwen2.5-VL-72B [39] served as the CoT generator for producing MLLM-generated CoTs. Progressive training began with supervised fine-tuning on binary and N-ary relation detection tasks for 2 epochs using template-based CoT, followed by fine-tuning on 4k MLLM-generated CoT samples. After that, GRPO RL training was conducted: 2.4k steps on the N-ary relation detection task, 3.6k steps on the binary relation detection task, and finally 2k steps on both tasks.
Appendix B Cognitive CoT Generation Prompt
Figure S1: Scene graph caption CoT generation prompt.
Figure S2: Grounded situation frame CoT generation prompt.
As mentioned in §3.2, we employ two distinct prompts (Figure S1 and Figure S2) to guide cognitive CoT generation for binary and N-ary relation detection, respectively. Each prompt strategically comprises: 1) a concise task definition outlining the detection objective; 2) an illustrative in-context example demonstrating the desired reasoning steps and output format; and 3) a specific instruction prompting the model to generate CoT of the current image.
Appendix C Scene Graph Format
Figure S3: Scene graph format prompt used for training.
The scene graph format prompt employed for GRPO training is displayed in Figure S3. Furthermore, to enable the model’s generation of think process, we utilize templated-CoT (c.f. §3.2) for supervised fine-tuning.
Appendix D Discussions

### v3

@@ Introduction
[…]
Evaluation Metrics. For binary relation detection, we followed the standard SGG evaluation protocol (tang2020unbiased; wang2024all) 1) Recall: The ratio of correctly predicted triplets to the total number of ground-truth triplets, with both label accuracy and bounding box IoU >= 0.5 for subject and object. 2) mean Recall (mRecall): The average recall across all predicate classes. For N-ary relation detection, we adopted the same evaluation metrics as (pratt2020grounded): 1) verb: the accuracy of verb prediction. 2) value: the noun accuracy for individual semantic role. 3) value-all (val-all): Overall noun prediction accuracy across all semantic roles in an event. 4) grounded-value (grnd): Bounding box accuracy for each semantic role, requiring an IoU >= 0.5 between predicted and ground-truth boxes. 5) grounded-value-all (grnd-all): entire bounding box accuracy across all semantic roles in an event.
Implementation Details. Refer to the Appendix.
@@
ISL (pratt2020grounded) $_{\text{ECCV'20}}$ Close-ended 39.36 30.09 18.62 22.73 7.72
JSL (pratt2020grounded) $_{\text{ECCV'20}}$ 39.94 31.44 18.87 24.86 9.66
GSRTR (cho2021grounded) $_{\text{BMVC'21}}$ 40.63 32.15 19.28 25.49 10.10
CoFormer (cho2022collaborative) $_{\text{CVPR'22}}$ 44.66 35.98 22.22 29.05 12.21
SituFormer (wei2021rethinking) $_{\text{AAAI'22}}$ 44.20 35.24 21.86 29.22 13.41
GSRFormer (cheng2022gsrformer) $_{\text{ACM MM'22}}$ 46.53 37 .48 23.32 31.53 14.23
OpenSU (liu2023opensu) $_{\text{ICCVW'23}}$ Open-ended 50.10 41.20 26.56 34.27 15.70
@@
Table 2: Performance(%) comparison with various GSR models on the SWiG (pratt2020grounded) dataset.
  Binary Relation NN-ary Relation
@@
Table 3: Performance comparison (%) across various cognitive CoT strategies trained on single binary relation detection and NN-ary relation detection, separately. Blue indicates metrics without correct verb constraints.
Figure 3: The statistics of binary and NN-ary relation rewards.
Comparison with State-of-the-Art Methods
We compared Relation-R1 for unified relation comprehension with the existing state-of-the-art binary relation detection (Table 1) and N-ary relation detection approaches in (Table 2).
[…]
Figure 4: Qualitative results of binary relation detection.
[…]
Diagnostic Experiment
@@
Reward Analysis. We visualized the reward variation curves for binary and N-ary relation detection in Figure 3, respectively. It can be seen that at the beginning stage, both tasks obtains decent rewards due to the foundation laid by SFT. With the gradual progress of GRPO, the rewards of both tasks show a gradually increasing trend. In binary relation detection, guided with progressive CoT yields a slightly enhanced overall reward trajectory, although the improvement is not significant. This might be since the gap between the CoT generated by MLLM (e.g., Qwen (Qwen2.5-VL)) and that produced by the template is not obvious. Conversely, N-ary relation detection exhibits a significant surge in reward, primarily attributed to the model’s emergent capability to generate synonymous grounded situation frames.
@@
Figure 5: Completion lengths of binary and NN-ary relation.
@@
The statistic of completion length during GRPO training is in Figure 5. Consistent with the reward trends, the completion lengths for binary relation detection exhibit a gradual increase as GRPO progresses, suggesting the generation of more detailed and descriptive outputs over time. Notably, the template-based CoT strategy yields significantly shorter completions compared to the more flexible MLLM-generated and progressive CoT strategies, since only the cognitive process is involved. The progressive CoT guidance, by enabling the model to explore synonymous relational expressions and incorporate more nuanced details, leads to an increase in the average completion length.
Qualitative Results. Figure 4 shows a qualitative example of binary relation detection. As seen, the visualized thinking process reveals how Relation-R1 effectively decomposes the complex task into constituent cognitive steps: object classification, localization, and relation inference. For instance, after identifying the central bed, Relation-R1 pinpoints related objects e.g., rug and chair, and explicitly grounds the inferred spatial relations (on, beside) between them. Our approach demonstrates the successful decomposition of the scene into accurate, grounded object-relation pairs.
Conclusion
This paper reveals key limitations in MLLM visual relation understanding, particularly for N-ary relations and language-induced bias. We propose Relation-R1, a unified two-stage framework: build a reasoning foundation and structured outputs via cognitive CoT-guided SFT, followed by GRPO, promoting visual-semantic grounding and robust generalization. Besides, progressively applied cognitive CoT guidance, transitioning from template-based to MLLM-generated reasoning, further enhances generalization, especially for synonymous N-ary relations. Experiments on PSG and SWiG validate Relation-R1’s effectiveness in both binary and N-ary relation detection. Employing progressive CoT-guided SFT with R1-inspired refinement, Relation-R1 advances MLLMs towards more human-like relational reasoning.
Acknowledgments
This work was supported by the National Natural Science Foundation of China Young Scholar Fund (62402408), the Hong Kong SAR RGC General Research Fund under Grant (16208823), and the Hong Kong SAR RGC Early Career Scheme (26208924). This research was partially conducted by ACCESS – AI Chip Center for Emerging Smart Systems, supported by the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government. This work was also supported by the Hong Kong SAR RGC General Research Fund under Grant (16208823) and the Hong Kong SAR RGC Early Career Scheme (26208924).
Summary of the Appendix
To enhance the understanding of the main paper, supplementary material containing further details are provided, organized as follows:
  • §A elaborates on the implementation details of Relation-R1.
  • §B introduces the cognitive CoT generation prompt.
  • §C details the scene graph format utilized in this work.
  • §D provides additional experimental results.
  • §E shows further qualitative results.
  • §F discusses limitations and broader impacts of Relation-R1.
Appendix A Implementation Details
We utilized Qwen2.5-VL-3B (Qwen2.5-VL) as our base model. Relation-R1 was trained on an 8xA100 80GB GPU server. During the SFT phase, we employed a total batch size of 32, with 4 samples processed per training step. For the GRPO component, we used a total batch size of 48. The learning rates were set to 2e-5 for SFT and 1e-6 for GRPO, respectively. Qwen2.5-VL-72B (Qwen2.5-VL) served as the CoT generator for producing MLLM-generated CoTs. The weighting hyperparameters $\alpha$ and $\beta$ were all set to 0.5 in the reward function. Progressive training began with supervised fine-tuning on binary and N-ary relation detection tasks for 2 epochs using template-based CoT, followed by fine-tuning on 4k MLLM-generated CoT samples. After that, GRPO RL training was conducted: 2.4k steps on the N-ary relation detection task, 3.6k steps on the binary relation detection task, and finally 2k steps on both tasks.
Appendix B Cognitive CoT Generation Prompt
Figure S1: Scene graph caption CoT generation prompt.
Figure S2: Grounded situation frame CoT generation prompt.
As mentioned above, we employ two distinct prompts (Figure S1 and Figure S2) to guide cognitive CoT generation for binary and N-ary relation detection, respectively. Each prompt strategically comprises: 1) a concise task definition outlining the detection objective; 2) an illustrative in-context example demonstrating the desired reasoning steps and output format; and 3) a specific instruction prompting the model to generate CoT of the current image.
Appendix C Scene Graph Format
Figure S3: Scene graph format prompt used for training.
The scene graph format prompt employed for GRPO training is displayed in Figure S3. Furthermore, to enable the model’s generation of the think process, we utilize template-based CoT for supervised fine-tuning.
Appendix D More Experimental Results
Table S1: Few-Shot GSR performance comparison.
@@ Appendix D More Experimental Results
GSRTR (cho2021grounded) $_{\text{BMVC'21}}$ One-Shot 0.35 0.10 0.01 0.08 0.00
CoFormer (cho2022collaborative) $_{\text{CVPR'22}}$ 0.63 0.23 0.02 0.14 0.01
@@
ww/$o$ RL 27.31 10.21 1.43 9.35 1.22
GSRTR (cho2021grounded) $_{\text{BMVC'21}}$ 4-Shot 6.95 4.32 1.83 2.79 0.65
CoFormer (cho2022collaborative) $_{\text{CVPR'22}}$ 12.39 8.27 3.93 5.28 1.25
@@
ww/$o$ RL 34.88 17.14 3.94 15.63 3.38
Few-Shot N-ary Relation Detection. We also evaluated Relation-R1’s few-shot generalization capability (Table S1). In the one-shot setting, Relation-R1 achieved dramatic improvements over prior work, reaching 31.67% (Verb) and 3.41% (Grnd-all) compared to CoFormer’s (cho2022collaborative) 0.63% and 0.01%, respectively. This advantage is maintained in the 4-shot setting, where Relation-R1 reached 38.02% (Verb) and 5.92% (Grnd-all), significantly surpassing previous methods. These results clearly demonstrate Relation-R1’s superior ability to generalize and perform robust N-ary relation detection, setting a new SOTA for few-shot GSR.
Joint Binary & N-ary Relation Detection.
Table S2: Performance comparison (%) across various cognitive CoT strategies trained on joint binary relation detection and NN-ary relation detection, separately. Blue indicates metrics without correct verb constraints.
  Binary Relation NN-ary Relation
@@
Beyond solely training Relation-R1 on binary or N-ary relation detection tasks, we also investigated the joint training of both tasks within the same Relation-R1 model. The results of this joint training are presented in Table S2333The Grnd-all metric excludes ground-truth bounding boxes represented as [-1, -1, -1, -1] for calculation.. From the analysis of these results, three primary conclusions can be drawn: 1) When utilizing both the SFT and RL strategies, our Relation-R1 consistently achieves higher performance across nearly all metrics for both binary and N-ary relation detection tasks. Similar to single-task training, the MLLM-generated CoT strategy underperforms in the N-ary task. This is likely due to the inherent generality of the CoT generated by Qwen (Qwen2.5-VL), which may not precisely align with the ground-truth annotations, consequently leading to lower Verb accuracy. 2) Our proposed progressive CoT-guided paradigm consistently achieves the highest performance across all evaluated metrics for both tasks in the joint learning setting. 3) The joint progressive CoT-guided training across both binary and N-ary tasks fosters a synergistic mutual promotion effect. This is evidenced by the notable gains observed: for instance, a Recall of 27.19% in Table S2 vs. 22.57% in Table 3 (for binary relation detection), and a Verb accuracy of 73.25% in Table S2 vs. 71.04% in Table 3 (for N-ary relation detection).
Appendix E Qualitative Results
@@
Figure S4: Performance(%) comparison between ASMv2 and Relation-R1 over all predicates on test set of PSG (yang2022panoptic) dataset. The orange area denotes the predicate distribution of the training set.
Figure S5: Qualitative results of binary relation detection.
Figure S6: Qualitative results of N-ary relation detection.
Figure S7: Qualitative results of synonymous in N-ary relation detection.
@@
Comparison over All Predicates. We visualized predicate-level comparison against ASMv2 (wang2024all) and the predicate distribution in Figure S4. As seen, Relation-R1 outperforms ASMv2 by a clear margin on both prevalent and rare predicates. These results clearly validated the effectiveness and efficiency of our approach for the SGG task in the open-ended setting.
@@
Binary and N-ary Relation Detection Results. We visualized some examples of binary and N-ary relation detection in Figure S5 and Figure S6, respectively. As seen, for binary relation detection (c.f. Figure S5), the visualized thinking process reveals how Relation-R1 effectively decomposes the complex task into constituent cognitive steps: object classification, localization, and relation inference. For instance, Relation-R1 can first focus on the main objects (e.g., “cat” and “train”) and then infer their positions and relationships. Similarly, for N-ary relation detection (c.f. Figure S6), Relation-R1 showcases its understanding by first identifying the core action (crowning) and then methodically grounding each associated semantic role – Agent (woman), Object (tiara), and Recipient (woman) – to its corresponding entity and bounding box. This structured analysis underscores Relation-R1’s ability to parse complex events involving multiple participants and their distinct roles within the scene.
@@
Synonymous Visualization. Figure S7 displays synonymous results observed in N-ary relation detection. As evinced by these results, our progressive CoT-guided training paradigm empowers the model to generate diverse synonymous expressions, encompassing entities, semantic roles, and activities. For instance, in Figure S7(top), Relation-R1 demonstrates the ability to output synonymous agents such as “person” and “man”. Furthermore, the model can generate semantically similar and plausible activities, exemplified by “barbecues” and “grills” (c.f. Figure S7(bottom)). Even within the same activity “rests”, it can differentiate and utilize varying semantic role names like “item” and “bodypart” (c.f. Figure S7(row 2)). These synonymous outputs indicate a more flexible and robust semantic understanding beyond rigid lexical matching, thereby significantly enhancing the model’s generalization capabilities to diverse linguistic formulations and real-world variations.
Appendix F Discussions